# Efficient Retrieval with Learned Similarities

### v2

@@ 3 Retrieval Algorithms
Gap of approximate top $K$: Let $q$ be a query and $X_{K}$ be the set of exact top $K$ items for the query $q$ from a set of items $X$. Let $X^{*}$ be the approximate top $K$ results, where $X^{*}\subseteq X$. Let $S=\min\{\phi(q,x),x\in X^{*}\}$ and $S^{\prime}=\max\{\phi(q,x),x\in X_{K}\setminus X^{*}\}$. We call $S_{\Delta}=S^{\prime}-S$ the gap of the top $K$ with $X^{*}$.
3.1 Exact algorithm
The brute-force algorithm to retrieve the exact top $K$ with MoL is to evaluate $\phi(q,x)$ for each query $q$ and item $x$. This algorithm can be prohibitively expensive if the number of items is large. Instead, we describe a more efficient two-pass algorithm to retrieve the exact top $K$ items as shown in Algorithm 1.
@@ 3.1 Exact algorithm
5:    $X_{p}\leftarrow\{g_{p}(x),x\in X\}$triangleright Can be preprocessed and materialized.
6:    $G\leftarrow G\cup TopKDotProduct(f_{p}(q),X_{p})$triangleright Retrieve top $K$ items for each pair of embeddings
@@
13:    $G^{\prime}\leftarrow G^{\prime}\cup RangeDotProduct(f_{p}(q),S_{min},X_{p})$$\triangleright$ Retrieve all items with a dot product higher than or equal to $S_{min}$ for $f_{p}(q)$ from $X_{p}$
14:return $BruteForceTopKMoL(q,G^{\prime})$$\triangleright$ Retrieve the top $K$ items from $G^{\prime}$ with MoL
We start by retrieving the top $K$ items with the highest dot product scores for each group of embeddings as the initial candidate set $G$ (line 1-4). Then we evaluate the MoL scores of the items in $G$ and find the minimal learned similarity score $S_{min}$ (line 5-8). Next we retrieve all items within a distance of $S_{min}$ with the query $q$ as the candidate set $G^{\prime}$ (line 9-11). Finally, we evaluate the MoL scores of the items in $G^{\prime}$, and return the top $K$ items with the highest scores (line 12).
@@
3.2 Approximate algorithms
In the exact algorithm shown in Algorithm 1, we need to retrieve all the items with a dot product higher than or equal to a threshold. When the threshold is a loose filter of the item set, which may happen when the dot product scores are skewed, $G^{\prime}$ can be large, and the evaluation of MoL over a large set of candidate items can still be expensive. Here, we describe two heuristics to approximately retrieve the top $K$ items and analyze their gap against the exact top $K$ algorithm.
@@ 3.2 Approximate algorithms
Algorithm 2 Approximate top $K$ algorithms.
@@
4:    $G\leftarrow TopKCandidate(q,X,K^{\prime})$triangleright Retrieve the top $K^{\prime}$ candidate items
5:    return $BruteForceTopKMoL(q,G,K)$$\triangleright$ Retrieve the top $K$ items with MoL
6:a query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the component-level embeddings $f_{p}(q),g_{p}(x)$
7:union of top $K$ items over $P$ component-level embeddings by dot product
@@
11:         $X_{p}\leftarrow\{g_{p}(x),x\in X\}$triangleright Can be preprocessed and materialized.
12:         $G\leftarrow G\cup TopKDotProduct(f_{p}(q),X_{p},K)$$\triangleright$ Retrieve the top $K$ items by dot product
13:    return $G$
14:a query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the component-level embedding $f_{p}(q),g_{p}(x)$
15:top $K$ items based on the averaged dot product, $\sum_{p}\langle f_{p}(q),g_{p}(x)\rangle/P$.
@@
18:    $X^{\prime}\leftarrow\{\sum_{p=1}^{P}g_{p}(x)/P,x\in X\}$triangleright Can be preprocessed and materialized.
@@
Here, we describe two heuristics to retrieve the candidate item set:
@@
Given a query $q$ and a set of items $X$, for each embedding set $p$, retrieve top $K$ items $X_{K,p}$ based on dot product ($\langle f_{p}(q),g_{p}(x)\rangle$). Return the union across $P$ queries.
The top $K$ per embedding heuristic returns the union of the top $K$ items for each embedding by dot product. We analyze the gap of this heuristic as follows:
@@
Upper bound of the gap of top $K$ per embedding: Let $X_{K,p}$ be the top $K$ items of the embedding set $p$ and $S=\max\{\phi(q,x),x\in X_{K+1,p}\}$. Let $S_{min}$ be the top $K^{th}$ MoL score of the items in $\cup_{p}X_{K,p}$, then the gap of $S_{\Delta}\leq S^{\prime}-S_{min}$.
Remark Note that there exists an MoL such that $S_{\Delta}=S-S_{min}$, i.e., when $\pi_{p}(q,x)=1$ for $x_{p}=\operatorname*{arg\,max}_{x,p}\{\langle f_{p}(q),g_{p}(x)\rangle,x\in X_{K+1,p}\setminus X_{K,p}\}$. Thus, the upper bound of $S_{\Delta}$ is tight.
@@
Given a query $q$ and a set of items $X$, return the top $K$ items with the highest average dot product $\sum_{p}\langle f_{p}(q),g_{p}(x)\rangle/P$.
@@
This heuristic is interesting for two reasons. First, the items retrieved by this heuristic are likely to be the top $K$ items of MoL when the weight distribution is more balanced. This complements the heuristic that retrieves top $K$ per embedding. Second, in the setup where the set of embedding pairs is constructed as the outer product of the embeddings of a query and those of an item (Equation 2), the average dot product can be efficiently preprocessed and materialized for the items, and the computation of the top $K$ average is then agnostic to the number of embedding sets $P=P_{q}\times P_{x}$.
Formally, let $P=P_{q}\cdot P_{x}$ be the number of embedding pairs, where $P_{q}$ is the number of embeddings of a query $q$ and $P_{x}$ is that of an item $x$. The average dot product can be computed as
$\frac{1}{P}\cdot\sum_{p=1}^{P}{\langle f_{p}(q),g_{p}(x)\rangle}=\frac{1}{P}\cdot\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}{\langle f_{p_{q}}(q),g_{p_{x}}(x)\rangle}=\frac{1}{P}\cdot\left\langle\sum_{p_{q}=1}^{P_{q}}f_{p_{q}}(q),\sum_{p_{x}=1}^{P_{x}}g_{p_{x}}(x)\right\rangle$ (3)
@@
Upper bound of the gap of combined top $K$. Let $X_{K,p}$ be the top $K$ items of the embedding set $p$ and $S_{min}$ as defined in Theorem 2. Let $X^{\prime}_{K}$ be the top $K$ items from top $K$ average. Let $S^{\prime}=\max\{\phi(q,x),x\in X\setminus(\cup_{p}X_{K,p}\cup X^{\prime}_{K})\}$. Then the gap of $S_{\Delta}\leq S^{\prime}-S_{min}$.
Remark Similar to Theorem 2, the upper bound of the gap is tight. In practice, we can configure the $K$ to be different for the two heuristics, i.e., $K_{1}$ and $K_{2}$. For example, when the weight distribution $\pi$ is more balanced, $K_{2}$ can be configured to be larger as the top $K$ average approach approximates MoL well while being more computationally efficient.
4 Evaluation
In this section, we evaluate the performance of the MoL based learned similarity and the efficiency of our retrieval algorithms discussed in Section 3.
4.1 Setup
Workloads.
We consider a classical retrieval task (next-item prediction) in recommendation systems [5, 45, 26, 42]. We employ three widely used datasets, the 1M and 20M subsets of MovieLens [13], and the largest Books subset of Amazon Reviews [29]. We utilize sequential user encoders, as they have been shown to achieve state-of-the-art results on these datasets [20, 42]. In the next-item prediction setting in sequential recommendations, the query $q$ corresponds to the user representation at a particular time step, which is then mapped to $P_{q}$ embeddings using an multi-layer perceptron (MLP) when using the MoL learned similarity, consistent with prior work. We discuss detailed hyperparameter settings in Appendix A. Table 2 shows the statistics of these three workloads.
@@ 4.1 Setup
Table 2: Workload statistics.
Model.
We implemented MoL on top of two popular sequential architectures, SASRec [20] and HSTU [43], with cosine similarity as the default similarity function (SASRec / HSTU rows in Table 3). The models are are implemented in PyTorch and trained with 2x NVIDIA RTX 6000 Ada GPUs.
Method HR@1 HR@10 HR@50 HR@200 MRR
ML-1M SASRec .0610 .2818 .5470 .7540 .1352
SASRec + MoL .0710 .3034 .5655 .7680 .1448
HSTU .0750 .3332 .5956 .7824 .1579
HSTU + MoL .0837 .3460 .6048 .7874 .1673
ML-20M SASRec .0653 .2883 .5484 .7658 .1375
SASRec + MoL .0790 .3124 .5685 .7784 .1551
HSTU .0962 .3557 .6146 .8080 .1800
HSTU + MoL .1008 .3697 .6244 .8124 .1880
Books SASRec .0058 .0306 .0754 .1431 .0153
SASRec + MoL .0083 .0402 .0970 .1813 .0203
HSTU .0101 .0469 .1066 .1876 .0233
HSTU + MoL .0150 .0613 .1292 .2167 .0315
Table 3: Evaluation of recommendation retrieval performance, utilizing the next-item prediction task.
@@
We use Recall (Hit Rate) as the main metric. It’s worth noting that HSTU + MoL sets new state-of-the-art results on MovieLens and Amazon Reviews (Table 3). Mixture-of-Logits (+ MoL rows) also consistently outperform dot products by an average of 24.2% in HR@1, 14.3% in HR@10, and 16.3% in MRR across the six scenarios, validating that learned similarities, in particular MoL, are not only theoretically expressive but also practically learnable. We evaluate our proposed retrieval algorithms relative to the best performing rows, “HSTU + MoL”, which exhaustively evaluates $\phi(q,x)$ over all $x$s and corresponds to BruteForce rows in the next section.
4.2 Top $K$ retrieval performance

### v4

@@ 3. Retrieval Algorithms
Gap of approximate top $K$: Let $q$ be a query and $X_{K}$ be the set of exact top $K$ items for the query $q$ from a set of items $X$. Let $X^{*}$ be the set of approximate top $K$ results, where $X^{*}\subseteq X$. Let $S=\min\{\phi(q,x),x\in X^{*}\}$ and $S^{\prime}=\max\{\phi(q,x),x\in X_{K}\setminus X^{*}\}$. We call $S_{\Delta}=S^{\prime}-S$ the gap of the top $K$ with $X^{*}$.
3.1. Exact algorithm
The brute-force algorithm to retrieve the exact top $K$ with MoL is to evaluate $\phi(q,x)$ for each query $q$ and item $x$. This algorithm can be prohibitively expensive if the number of items is large. Instead, we describe a more efficient two-pass algorithm to retrieve the exact top $K$ items as shown in Algorithm 1 (example in Appendix B).
@@ 3.1. Exact algorithm
5:    $X_{p}\leftarrow\{g_{p}(x),x\in X\}$triangleright Can be preprocessed.
6:    $G\leftarrow G\cup TopKDotProduct(f_{p}(q),X_{p})$triangleright Retrieve top $K$ items for each pair of embeddings.
@@
13:    $G^{\prime}\leftarrow G^{\prime}\cup RangeDotProduct(f_{p}(q),S_{min},X_{p})$$\triangleright$ Retrieve all items $x\in X_{P}$ with $\langle f_{p}(q),x\rangle\geq S_{min}$.
14:return $BruteForceTopKMoL(q,G^{\prime})$triangleright Retrieve the top $K$ items from $G^{\prime}$ with MoL.
We start by retrieving the top $K$ items with the highest dot product scores for each group of embeddings as the initial candidate set $G$ (line 1-4). Then we evaluate the MoL scores of the items in $G$ and find the minimal MoL score $S_{min}$ (line 5-8). Next we retrieve all items within a distance of $S_{min}$ with the query $q$ as the candidate set $G^{\prime}$ (line 9-11). Finally, we evaluate the MoL scores of the items in $G^{\prime}$, and return the top $K$ items with the highest scores (line 12).
@@
3.2. Approximate algorithms
In the exact algorithm shown in Algorithm 1, we need to retrieve all the items with a dot product higher than or equal to a threshold. When the threshold is a loose filter of the item set, which may happen when the dot products are skewed, $G^{\prime}$ can be large, and the evaluation of MoL over a large number of candidates can be expensive. Here, we describe two heuristics to approximately retrieve the top $K$ items and analyze their gap against the exact algorithm.
@@ 3.2. Approximate algorithms
Algorithm 2 Approximate top-$k$ algorithms.
@@
4:    $G\leftarrow TopKCandidate(q,X,K^{\prime})$triangleright Retrieve the top $K^{\prime}$ candidates.
5:    return $BruteForceTopKMoL(q,G,K)$$\triangleright$ Retrieve the top $K$ items from $G$ with MoL.
6:a query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the $P$ component-level embedding pairs $f_{p}(q),g_{p}(x)$
7:union of top $K$ items over $P$ embedding pairs by dot product
@@
11:         $X_{p}\leftarrow\{g_{p}(x),x\in X\}$triangleright Can be preprocessed.
12:         $G\leftarrow G\cup TopKDotProduct(f_{p}(q),X_{p},K)$$\triangleright$ Retrieve the top $K$ items by dot product for the $p$-th embedding pair.
13:    return $G$triangleright Dedup’ed set of the top $K$ item for each of the $P$ queries
14:a query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the component-level embeddings $f_{p}(q),g_{p}(x)$
15:top $K$ items with averaged dot product, $\sum_{p}\langle f_{p}(q),g_{p}(x)\rangle/P$
@@
18:    $X^{\prime}\leftarrow\{\sum_{p=1}^{P}g_{p}(x)/P,x\in X\}$triangleright Can be preprocessed.
@@
Here, we describe two heuristics to retrieve the candidate items:
@@
Given a query $q$ and a set of items $X$, for each of the $p$ embedding pairs, retrieve top $K$ items $X_{K,p}$ based on dot product ($\langle f_{p}(q),g_{p}(x)\rangle$). Return the union across $P$ queries.
The top $K$ per embedding heuristic returns the union of the top $K$ items for each embedding pair (group) by dot product. We analyze the gap of this approach as follows:
@@
Upper bound of the gap of top $K$ per embedding: Let $X_{K,p}$ be the top $K$ items of the embedding set $p$ and $S=\max\{\langle f_{p}(q),g_{p}(x)\rangle,x\in X_{K+1,p}\setminus X_{K,p},\forall p\}$. Let $S_{K}$ be the $K^{th}$ largest MoL score of the items in $\cup_{p}X_{K,p}$, then we have $S_{\Delta}\leq S-S_{K}$.
Remark This gap (error bound) bounds the maximal difference between the $K^{th}$ largest MoL score from the set of items retrieved by the heuristic and the actual $K^{th}$ largest MoL score. In addition, any retrieved item $x$ with $\phi(q,x)\geq S$ belongs to the actual top $K$ items. Note that there exists an MoL such that $S_{\Delta}=S-S_{K}$, i.e., when $\pi_{p}(q,x)=1$ for $x,p=\operatorname*{arg\,max}_{x,p}\{\langle f_{p}(q),g_{p}(x)\rangle,x\in X_{K+1,p}\setminus X_{K,p},\forall p\}$. Thus, the upper bound of $S_{\Delta}$ is tight. In practice, we can calculate a looser bound with the items retrieved by per embedding top $K$, i.e., from $X_{K,p}$. We provide an example in Appendix B.
@@
Given a query $q$ and a set of items $X$, return the top $K$ items with the highest average dot product defined as $\sum_{p}\langle f_{p}(q),g_{p}(x)\rangle/P$.
@@
This heuristic is interesting for two reasons. First, the items retrieved by this heuristic are likely to be the top $K$ items of MoL when the weight distribution is more balanced. This complements the heuristic that retrieves top $K$ per embedding. Second, in the setup where the set of embedding pairs is constructed as the outer product of the embeddings of a query and those of an item (Equation 2), the average dot product can be efficiently preprocessed and materialized for the items, and the computation of the top $K$ average is then agnostic to the number of embedding pairs.
Formally, let $P=P_{q}\times P_{x}$ be the number of embedding pairs, where $P_{q}$ is the number of embeddings of a query $q$ and $P_{x}$ is that of an item $x$. The average dot product can be computed as
(5) $\frac{1}{P}\cdot\sum_{p=1}^{P}{\langle f_{p}(q),g_{p}(x)\rangle}$ $=$ $\frac{1}{P}\cdot\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}{\langle f_{p_{q}}(q),g_{p_{x}}(x)\rangle}$
(6) $=$ $\frac{1}{P}\cdot\left\langle\sum_{p_{q}=1}^{P_{q}}f_{p_{q}}(q),\sum_{p_{x}=1}^{P_{x}}g_{p_{x}}(x)\right\rangle$
@@
Upper bound of the gap of combined top $K$. Let $X_{K,p}$ be the top $K$ items of the embedding set $p$ and $S_{K}$ as defined in Theorem 3. Let $X^{\prime}_{K}$ be the top $K$ items from top $K$ average. Let $S^{\prime}=\max\{\langle f_{p}(q),g_{p}(x)\rangle,x\in X\setminus(X_{K,p}\cup X^{\prime}_{K}),\forall p\}$. Then the gap $S_{\Delta}$ satisfies $S_{\Delta}\leq S^{\prime}-S_{K}$.
Remark Similar to Theorem 3, the upper bound of this gap is tight. In practice, we can configure the $K$ to be different for the two heuristics, i.e., $K_{1}$ and $K_{2}$. For example, when the weight distribution $\pi$ is more balanced, $K_{2}$ can be configured to be larger as the top $K$ average approach approximates MoL well while being more computationally efficient.
4. Evaluation
In this section, we evaluate the performance of MoL based learned similarity with the proposed load balancing loss, and the efficiency of retrieval algorithms discussed in Section 3. Our code and model checkpoints are available at https://github.com/bailuding/rails.
4.1. Workloads
We benchmark MoL with the proposed load balancing loss $\mathcal{L}_{MI}$, on top of state-of-the-art baselines in recommendation systems and question answering. We describe workloads used below.
Recommendation Systems.
We consider three widely used datasets, the 1M and 20M subsets of MovieLens (Harper and Konstan, 2015), and the largest Books subset of Amazon Reviews (McAuley et al., 2015). Sequential retrieval models have been shown to achieve state-of-the-art results on these datasets (Hidasi et al., 2016; Kang and McAuley, 2018; Zhai et al., 2024). In these settings, sequential encoders, like RNNs or Transformers, are used to map user representations at time $t$ – e.g., in a commonly used setting shown in Figure 2, the list of items in user history up until time $t$, $\Phi_{0},\ldots,\Phi_{t}$ – to $\mathbb{R}^{d}$, and the model is trained to autoregressively predict the next item $x_{t+1}$. We hence compare MoL with the proposed regularization loss on top of two popular backbones used for sequential retrieval models, SASRec (Kang and McAuley, 2018) and HSTU (Zhai et al., 2024), against cosine similarity baselines. We utilize user id-based embeddings discussed in Section 2.2 and MLPs to parameterize the $P_{Q}$ query-side and the $P_{X}$ item-side features.
Question Answering (QA)
Natural Questions (NQ) (Kwiatkowski et al., 2019) is commonly used to evaluate state-of-the-art neural retrieval models, including dense retrieval (Karpukhin et al., 2020; Ni et al., 2022b) and generative retrieval (De Cao et al., 2021; Tay et al., 2022; Wang et al., 2022b; Sun et al., 2023) approaches in recent years. The most commonly used version (Tay et al., 2022; Wang et al., 2022b; Sun et al., 2023), which we reuse in our work, is often referred to as NQ320k. NQ320k consists of 320k query-items pairs, where the items are from Wikipedia pages and the queries are natural language questions. We utilize special aggregation tokens discussed in Section 2.2 to parameterize embeddings in MoL, and compare MoL with popular sparse retrieval methods (Robertson and Zaragoza, 2009; Nogueira and Lin, 2019), dense retrieval methods (Karpukhin et al., 2020; Ni et al., 2022a, b), and generative retrieval methods (De Cao et al., 2021; Tay et al., 2022; Bevilacqua et al., 2022; Wang et al., 2022b; Sun et al., 2023). Consistent with recent work (Zhuang et al., 2023; Wang et al., 2022b; Sun et al., 2023), we use the pre-trained query generation model from DocT5Query (Nogueira and Lin, 2019) to generate synthetic (query, item) pairs for data augmentation.
Table 2 summarizes the statistics of these four workloads.
@@ 4.1. Workloads
NQ320K 307,373 109,739 4 4 768
Table 2. Workload statistics.
4.2. Quality of MoL-based Learned Similarity
@@
We use Recall (Hit Rate) as the main metric. We report Hit Rate@{1, 10, 100} and Mean Reciprocal Rank (MRR) on NQ320K, following (Wang et al., 2022b; Sun et al., 2023), and Hit Rate@{1, 10, 50, 200} on ML-1M, ML-20M, and Books, following (Zhai et al., 2023, 2024).
Hyperparameter Settings.
We set the weight $\alpha$ for the proposed load balancing loss $\mathcal{L}_{MI}$ to $0.001$ for all experiments. We reuse baseline settings for most other hyperparameters, including learning rate, number of examples used for in-batch negative sampling, etc., with detailed discussions in Appendix A. For the NQ320K dataset, we reuse SEAL (Bevilacqua et al., 2022) and NCI (Wang et al., 2022b) results reported by (Wang et al., 2022b), and results for other models as reported by (Sun et al., 2023). The Sentence-T5 (Ni et al., 2022a), GENRE (De Cao et al., 2021), DSI (Tay et al., 2022), SEAL (Bevilacqua et al., 2022), DSI+QG (Zhuang et al., 2023), NCI (Wang et al., 2022b), and GenRet (Sun et al., 2023) rows are all finetuned from T5-base, consistent with MoL, to ensure a fair comparison. All other results are implemented in PyTorch, and are trained with 1x/2x 48GB GPUs for the recommendation datasets and 4x 80GB GPUs for the QA datasets.
Method HR@K MRR
K=1 K=10 K=50 K=200
ML-1M dataset
SASRec (Kang and McAuley, 2018) .0610 .2818 .5470 .7540 .1352
SASRec + MoL .0697 .3036 .5617 .7667 .1441
HSTU (Zhai et al., 2024) .0750 .3332 .5956 .7824 .1579
HSTU + MoL .0884 .3465 .6022 .7935 .1712
HSTU + MoL abl. $\mathcal{L}_{MI}$0847 .3417 .6011 .7942 .1662
ML-20M dataset
SASRec (Kang and McAuley, 2018) .0653 .2883 .5484 .7658 .1375
SASRec + MoL .0778 .3102 .5682 .7779 .1535
HSTU (Zhai et al., 2024) .0962 .3557 .6146 .8080 .1800
HSTU + MoL .1010 .3698 .6260 .8132 .1881
HSTU + MoL abl. $\mathcal{L}_{MI}$0994 .3670 .6241 .8128 .1866
Books dataset
SASRec (Kang and McAuley, 2018) .0058 .0306 .0754 .1431 .0153
SASRec + MoL .0095 .0429 .0915 .1635 .0212
HSTU (Zhai et al., 2024) .0101 .0469 .1066 .1876 .0233
HSTU + MoL .0156 .0631 .1308 .2173 .0324
HSTU + MoL abl. $\mathcal{L}_{MI}$0153 .0625 .1286 .2172 .0321
Table 3. Evaluation of performance for sequential retrieval models on MovieLens and Amazon Reviews.
Method HR@K MRR
K=1 K=10 K=100
Sparse retrieval
BM25 (Robertson and Zaragoza, 2009) .297 .603 .821 .402
DocT5Query (Nogueira and Lin, 2019) .380 .693 .861 .489
Dense retrieval
DPR (Karpukhin et al., 2020) .502 .777 .909 .599
Sentence-T5 (Ni et al., 2022a) .536 .830 .938 .641
GTR-Base (Ni et al., 2022b) .560 .844 .937 .662
Generative retrieval
GENRE (De Cao et al., 2021) .552 .673 .754 .599
DSI (Tay et al., 2022) .552 .674 .780 .596
SEAL (Bevilacqua et al., 2022) .570 .800 .914 .655
DSI+QG (Zhuang et al., 2023) .631 .807 .880 .695
NCI (Wang et al., 2022b) .659 .852 .924 .731
GenRet (Sun et al., 2023) .681 .888 .952 .759
Learned similarities
MoL .685 .919 .970 .773
MoL abl. $\mathcal{L}_{MI}$673 .919 .968 .767
Table 4. Evaluation of performance for QA retrieval models finetuned from language models on Natural Questions.
Results.
Across the six recommendation scenarios utilizing different sequential encoder backbones, Mixture-of-Logits (MoL rows) consistently outperform the state-of-the-art dense retrieval baselines (dot products) by an average of 29.1% in HR@1, 16.3% in HR@10, and 18.1% in MRR (Table 3). On the widely used Natural Questions QA dataset, MoL outperforms all recent generative retrieval approaches as well as strong dense- and sparse- retrieval baselines (Table 4). These results validate that learned similarities, in particular MoL, are not only theoretically expressive but also practically learnable, improving retrieval quality across heterogeneous scenarios, including sequential retrieval models for Recommendations and finetuning LMs for Question Answering.
Ablation Studies.
We conduct ablation studies for the proposed mutual information-based load balancing loss relative to the best performing method for each dataset (“abl. $\mathcal{L}_{MI}$” rows). Results show that our proposed $\mathcal{L}_{MI}$ loss improves HR@1 by 2.4%, HR@10 by 0.8% and MRR by 1.4% across the four datasets. In particular, our proposed $\mathcal{L}_{MI}$ loss enables MoL to outperform the best generative retrieval approach on NQ320K, GenRet (Sun et al., 2023), across all metrics.
4.3. Top $K$ retrieval performance